Oxygen is present in a flask of 1.12 L capacity at a pressure of 7.6×10 -10 mm of Hg at0°c.The no of oxygen molecules in the flask is ( And:3×10-10

Grade:12

1 Answers

Arun
6 years ago
Dear Harsha
 
NOTE THIS STEP : First we should calculate the number of moles of the gas under the given condition by the relation PV = nRT
 
Here P = 7.6 * 10-10 mm Hg
 
= 7.6 *10-10/760 atm. = 1 * 10-12 atm.
 
V = 1.2 litre
 
T = 273 + 0 = 273K
 
R = 0.082 litre atm./K/mol
 
Putting the values in equation
 
n = PV/RT = 1 *10-12 *1.2/0.082 *273 moles
 
now since 1 mole = 6.023 * 1023 molecules
 
1.2*10-12/0.082 *273 moles =1.2* 6.023 *1023 *10-12/0.082 *273 molecules
 
= 3.24 * 1010 molecules
 
Regards
Arun (askIITians forum expert)